### v2

# Forecasting from Clinical Textual Time Series: Adaptations of the Encoder and Decoder Language Model Families

Abstract
Clinical case reports encode rich, temporal patient trajectories that are often underexploited by traditional machine learning methods relying on structured data. In this work, we introduce the forecasting problem from textual time series, where timestamped clinical findings–extracted via an LLM-assisted annotation pipeline–serve as the primary input for prediction. We systematically evaluate a diverse suite of models, including fine-tuned decoder-based large language models and encoder-based transformers, on tasks of event occurrence prediction, temporal ordering, and survival analysis. Our experiments reveal that encoder-based models consistently achieve higher F1 scores and superior temporal concordance for short- and long-horizon event forecasting, while fine-tuned masking approaches enhance ranking performance. In contrast, instruction-tuned decoder models demonstrate a relative advantage in survival analysis, especially in early prognosis settings. Our sensitivity analyses further demonstrate the importance of time ordering, which requires clinical time series construction, as compared to text ordering, the format of the text inputs that LLMs are classically trained on. This highlights the additional benefit that can be ascertained from time-ordered corpora, with implications for temporal tasks in the era of widespread LLM use.
**footnotetext: These authors contributed equally to this work
1 Introduction
Clinical risk forecasting is a cornerstone problem in health care decision making that guides eligibility and reasoning about interventions. Because risk prediction performs best with access to rich data streams, great lengths are taken to capture the relevant health factors that affect decision making. However, while many forecasting tasks are performed on structured tabular data, evidence suggests that, despite widespread efforts to collect large amount of data in structured format, only a small fraction of structured data overlaps with that documented in text (Seinen et al., 2025). Furthermore, empirically we observe that predictions are considerably better when text is integrated (Kline et al., 2022).
A complementary data stream in healthcare comes from the textual data that providers use in documenting their encounters. Likely the most comprehensive part are the case reports and discharge summaries, which are a sort of “post-mortem”, retrospective discussion of the relevant factors for decision making and observed progression. However, without further effort, we are unable to leverage this data stream for forecasting because the times that are documented are interwoven in the narrative passage.
To construct and learn from these patient trajectories, this work investigates the development of a forecasting pipeline for textual time series from retrospective analysis of clinical trajectories from case reports. Using the framework in Wang and Weiss (2025), we leverage extracted temporally located clinical findings, namely lists of (event, time) tuples that represent the patient encounter. The mapping from case report to time-localized events allows us to make anytime forecasts by masking those events that occur after the requested forecast time. Without this annotation process, it would be unclear which findings could be used so as to avoid “causal leakage” from events occurring in the future.
Our approach differs from other efforts in the NLP and clinical informatics community and is driven by our particular emphasis on high-performance forecasting and temporally-aware risk characterization. For example, our definition of clinical finding is related to clinical entity and clinical concept definitions from the field of named entity recognition. However, we avoid mapping to a discrete set of entities to increase the availability of contextual information for forecasting, and thus require an alternative annotation guideline (Uzuner et al., 2011). With respect to our temporal annotations, our approach differs from the classical temporal relations task, which, while capturing relative ordering and overlap of events, makes it difficult to perform risk calculations and multimodal alignments that are functions of time (Kohane, 1987; Zhou and Hripcsak, 2007; Sun et al., 2013).
Figure 1: Example case report (left) with temporal representations as timelines (middle) and text-ordered, event-time tuples (right).
We use LLMs to assist us in this task, so it is natural to question whether we need this extraction/annotation step at all, since a multitude of BERT-based and LLM models have consumed these, similar post-mortem reports, and more during training (Peng et al., 2020; Weber et al., 2024). From an architectural perspective, both encoder (BERT-style) and decoder (LLM-style) models will not perform the temporal masking necessary to encode leakage-free embeddings. For BERT models, which use full attention modules with masking, the probability that the random masking recovers a time-ordered representation is vanishingly small (tested empirically on the example timeline in Figure 1 (right) with default BERT masking probability of 0.15; probability of a time-ordered masking: 2/100,000). For decoder models, their causal attention mask means that embeddings are processed in text-order rather than time-order. Given that the concordance, the fraction of the time pairwise temporal ordering are correct, between text-ordered and time-ordered events is lower than 1 in most reports (c=0.23 in the example in Figure 1), we would expect suboptimal temporal reasoning capabilities.
As we will see, we empirically find that these models do not forecast well without further training. If we can address these textual-temporal discrepancies, we expect to be able to substantially improve LLM capabilities at clinical forecasting, which could lead to downstream improvements in clinical temporal reasoning, risk estimation, causal reasoning, and decision making.
1.1 Contributions
  1. 1.
    Annotation and Extraction Pipeline: Building on methods of Noroozizadeh and Weiss (2025) for extracting information from the PubMed Open Access corpus, we develop a pipeline that converts case reports into textual time series using both regex-based filtering and LLM-assisted extraction. The pipeline produces temporally localized (event, time) tuples, mitigating potential causal leakage by ensuring that only pre-cutoff events inform forecasts.
  2. 2.
    Comprehensive Model Comparison: We rigorously evaluate a broad suite of models across multiple forecasting tasks. Our study compares instruction-tuned decoder-based models (both via fine-tuning with MLP heads and prompting strategies) against encoder-based models with task-specific fine-tuning. Experiments on event occurrence prediction, temporal ordering, and survival analysis reveal complementary strengths: no single modeling approach consistently outperforms across all tasks, highlighting the importance of aligning model choice with the specific forecasting objective.
  3. 3.
    Analysis of Temporal versus Text Ordering: We perform a detailed investigation into the impact of ordering on performance by contrasting training on time-ordered versus text-ordered annotations. Our experiments demonstrate that preserving the clinical narrative’s natural order can enhance generalization on external datasets, while explicit time ordering can improve ranking performance under specific conditions.
  4. 4.
    Sensitivity Analysis of Temporal Masking: Through systematic dropout experiments, we analyze the effects of randomly masking portions of the clinical history on forecasting accuracy and event ordering. The findings reveal that while classification performance (F1 score) deteriorates with higher levels of masking, the concordance index remains largely stable—highlighting differing sensitivities of binary prediction and ranking tasks to historical context in our textual time series setup.
Generalizable Insights about Machine Learning in the Context of Healthcare
Extensive research has been in the clinical NLP community towards temporal understanding of clinical narratives (Kohane, 1987; Leeuwenberg and Moens, 2020; Cheng and Weiss, 2023). While much has focused on temporal relations, relative-time event tuples facilitate the use of longitudinal and survival models so long as they can use free-text inputs and outcomes (possibly via mappings to an ontology). Whereas the primary focus in this space is on the entity and temporal extraction tasks, downstream application studies remain infrequent. Our study seeks to close thie gap by applying forecasting methodologies to these data streams to characterize the performance of leading machine learning and NLP approaches. A better empirical understanding of the strength and weakness of these approaches will enable future work to advance improved textual time series methods. Our investigation also highlights that risk estimation and event anticipation is not optimally captured in LLMs (as we demonstrate in contrasts between fine-tuning and prompt-based strategies), suggesting the need for careful investigation of LLMs when applied to assessment of risk.
2 Related Work
Temporal Information Extraction from Clinical Text:
Extracting timelines from clinical narratives has been a longstanding challenge in biomedical NLP. Early efforts such as the i2b2 2012 challenge provided annotated datasets for temporal relation extraction, focusing on identifying and ordering events in discharge summaries (Uzuner et al., 2011). Follow-up work constructed timelines from these narratives by linking clinical events to timestamps or temporal expressions (Leeuwenberg and Moens, 2020; Frattallone-Llado et al., 2024). These methods often relied on pre-defined event spans and excerpts, rather than full-document processing. In contrast, following Noroozizadeh and Weiss (2025), we assign timepoints to clinical findings from full-length case reports, allowing for finer temporal resolution. While some works integrate structured EHR data to anchor timelines, our approach focuses on timeline construction and forecasting using text alone, which is essential for corpora such as the PubMed Open-Access Subset, where no structured metadata is available. Our annotation strategy enables direct supervision of event-time alignment, overcoming the limitations of earlier span-only annotations. This aligns with the broader need in sepsis and critical care research to incorporate temporality into clinical phenotyping (Johnson et al., 2018; Kamran et al., 2024; Seinen et al., 2025), especially given the high missingness in structured data (Johnson et al., 2016, 2023).
Predictive Modeling with Clinical Text:
Incorporating unstructured text (clinical notes, narratives) into predictive models has been explored in various contexts, primarily for outcomes like mortality or readmission. Notably, ClinicalBERT and related transformer models pre-trained on medical notes have achieved state-of-the-art results in predicting hospital readmission and mortality (Huang et al., 2019; Gu et al., 2021). These works highlight that text contains complementary information to structured data, for example, documenting patient symptoms or social factors not in coded form. However, directly modeling text is challenging due to noise and length; researchers often rely on the full text or bag-of-words representations, which lose temporal granularity. Our approach differs by preserving the chronological sequence of events mentioned in text, effectively treating narrative data as a time series. This connects to recent studies on re-purposing LLMs for traditional time-series forecasting tasks: techniques like Time-LLM have shown that language models can be “re-programmed” to handle numeric sequences by mapping time points to token sequences (Jin et al., 2023). In the clinical domain, our work is the first to apply language models to text-derived time series, bridging free-text analysis with time-aware prediction. We also relate to prior sepsis prediction models. Traditional models use vital sign trends or clinical scores (e.g., SOFA, SAPS) to predict sepsis onset or mortality (Hou et al., 2020; Noroozizadeh et al., 2023). Unlike these, our method relies solely on narrative descriptions rather than real-time vitals, and is thus complementary: it demonstrates how knowledge embedded in case reports (which often include the clinicians’ interpretations and context) can be harnessed for prediction. This could be especially useful in settings where detailed narratives are available (such as case report literature or perhaps patient histories) but structured data is sparse.
Large Language Models in Healthcare:
The advent of powerful LLMs like GPT-3 and GPT-4 has sparked interest in various healthcare applications, from medical question-answering to clinical note summarization. These models, trained on massive text corpora, contain a wealth of medical knowledge and the ability to perform reasoning to some extent. For instance, GPT-4 has been tested on medical board exam questions with impressive results, and pilot studies show LLMs can abstract patient records for quality reporting (Boussina et al., 2024). Nevertheless, using LLMs for forecasting in healthcare is relatively uncharted territory. A limitation of out-of-the-box LLMs is their tendency to generate plausible-sounding but unverified content (hallucinations) and their lack of alignment with specific clinical data distributions. Our benchmark explicitly evaluates LLMs through prompting (zero-shot and few-shot) to see if they can ”forecast” events when given a patient’s story so far. We also compare that against fine-tuned models that are directly trained on the task, to quantify the gap between general-purpose and task-specific modeling in this context. This comparison adds to the growing body of evidence that, while large models have broad knowledge, domain-specific fine-tuning or structuring of input is often necessary for reliable performance in clinical tasks (Huang et al., 2019; Gu et al., 2021). Some recent analyses suggest that medical-domain fine-tuning may not consistently improve performance over foundation models (Jeong et al., 2024), further motivating the need for task- and input-specific evaluation setups like ours.
3 Methods
 [^1] We will be releasing our code upon publication.
We outline the task specification, define key terms, describe the extraction and annotation process, and present the evaluation criteria (Figure 2).
Figure 2: Flowchart of the T2S2 forecasting analysis pipeline
3.1 Dataset extraction
We utilized the PubMed Open Access (PMOA) Subset (as of April 25, 2024) for our analysis. Building on the method described in Wang and Weiss (2025), we extract the body of each case by selecting the text between the markers ==== Body and ==== Ref, which serve as section delimiters in the PMOA corpus. To maintain relevance, we include only those documents that contain case-insensitive matches to the regular expressions (case report|case presenta) and year-? ?old. Prior work has shown that this regex-based approach demonstrated notably higher specificity compared to PubMed metadata on a random sample of 20 case reports Noroozizadeh and Weiss (2025).
As demonstrated in prior work Noroozizadeh and Weiss (2025), we applied additional filtering to identify potential sepsis-related case reports using the regular expressions sepsi|septic and critical|intensive care. We then used LLM-based queries to extract key attributes, including sepsis diagnosis, patient count, age, and gender. For this step, both GPT-4o and Llama-3.1-8B-Instruct were employed. Reports describing more than one case were excluded. A case report was retained if either model classified it as involving a Sepsis-3 diagnosis.
This multi-step process yielded a total of 2319 Sepsis-3 case reports. From this set, we randomly selected two subsets for evaluation: a group of 10 reports (sepsis-10), which underwent expert clinical annotation to serve as a gold standard, and a larger group of 100 reports (sepsis-100) used for broader testing (Figure 2).
3.2 Textual time-series annotation
As established in previous work Noroozizadeh and Weiss (2025), we aimed to construct a textual time series case report corpus of Sepsis-3 patients from the PMOA sepsis case reports. A textual time series refers to a sequence of clinical findings, each paired with a timestamp (either absolute or relative to the case presentation time), corresponding to an individual patient. Here, a clinical finding is defined as a free-text expression describing an entity that pertains to or may impact a person’s health.
A textual time-series is different from conventional approaches to clinical concept annotation (e.g., Sun et al. (2013); Uzuner et al. (2011)) by extending the span of extracted text to better capture the specificity and contextual meaning of each clinical finding. In our work, the interpretation of a clinical finding differs from the i2b2 concept guidelines in the following respects : (1) Clinical findings are not restricted to single prepositional phrases following a markable entity. For instance, instead of splitting “pain in chest that radiates substernally” into less informative parts, we retain the full phrase as a single, complete finding to preserve its meaning and context. (2) To improve clarity, coordinated phrases are broken into individual findings. For example, “metastases in the liver and pancreas” becomes “metastasis in the liver” and “metastasis in the pancreas.”
We used DeepSeek-R1-UD-IQ1 and Llama-3.3-70B-Instruct models to generate clinical textual time series from the 2319 sepsis case reports. We refer to this as Textual Time-Series for Sepis (T2S2) The exact LLM extraction prompt can be found in Appendix E. Our dataset was split into training (n = 1795), testing T2S2-test (n = 244) and the 2 external validation sets sepsis 10 (n = 10) and sepsis 100 (n = 90). The training set was further split into 80% for model training and 20% for model validation (e.g. early stopping).
3.3 Task Definition
Using the timelines extracted from the annotations, we defined two primary forecasting tasks and a survival analysis task:
Event Occurrence Prediction:
Given a prefix of the clinical timeline (all events up to a certain time $t$), the model is tasked with predicting whether each of the immediate next $k$ events occurs within a specified time horizon. This setup is repeated across multiple time cut-offs to simulate an “online” forecasting scenario, where the model must output a binary label for each of the next $k$ events: does this event occur within $h$ hours after $t$? Time horizons used include 1 hour, 24 hours (1 day), and 168 hours (1 week). The task is framed as a series of binary classification problems, with one binary decision per event. Evaluation is based on precision, recall, and F1 score, averaged across event positions for each time horizon.
Temporal Ordering Prediction:
This task assesses the model’s ability to reconstruct the correct sequence of future events. At each time cut-off $t$, we extract the next $k$ events from the timeline and remove their timestamps. The model must correctly output a permutation over these events that matches their true chronological order. This is framed as a ranking task, evaluated by computing the pairwise concordance between the predicted and true orderings (e.g., proportion of correctly ordered pairs). This tests whether the model can infer temporal progression from unordered event content alone.
Survival Analysis for Mortality Time:
We include a classical survival analysis task to model time until death. Many case reports specify whether the patient died and, if so, when (e.g., “the patient died on hospital day 10”), which enables us to define a time-to-event outcome. For this task, we evaluate models at predefined cut-off times – specifically at 0 hours (admission), 24 hours (1 day), and 168 hours (1 week) – and use the extracted event history up to each cut-off as input. A survival model is trained to predict the probability of survival over time beyond each cut-off. We evaluate performance using the time-dependent concordance index to assess how well the predicted survival times align with actual outcomes.
3.4 Modeling Approaches
To evaluate performance on the event forecasting and survival prediction tasks, we implement three modeling paradigms: (i) large language models (LLMs) with fine-tuned heads, (ii) prompted LLMs without gradient updates (zero- or few-shot), and (iii) BERT-style encoder models with task-specific fine-tuned heads. In Appendix C and  D, we provide detailed explanations of how we approach the event occurrence prediction, temporal ordering prediction, and survival analysis tasks using encoder-only models.
Fine-tuned LLMs:
We apply instruction-tuned decoder-only models from the LLaMA and DeepSeek families: Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, and their distilled variants. Each model is paired with a lightweight multilayer perceptron (MLP) head trained for classification or ranking, depending on the task. The input consists of a text-formatted prefix of events (e.g., a clinical timeline up to time $t$), optionally accompanied by an instruction template. The output layer produces task-specific predictions: binary labels for event occurrence or a permutation over $k$ events for ordering. Training is conducted using cross-entropy loss for classification and pairwise ranking loss for ordering.
Prompted LLMs:
In the zero- or few-shot setting, we use the same LLM architectures as above, but without any fine-tuning. Instead, we supply structured prompts at inference time to guide the model toward task-specific outputs. Each prompt includes: (1) a system instruction establishing the model’s role (e.g., ”You are an expert physician.”), (2) a user instruction describing the prediction task, and (3) one or more input-output examples in a few-shot format. Prompt designs are customized for each task and are detailed in Appendix F. Model outputs are parsed to extract binary labels or ordered lists, and errors in generation or ambiguity are excluded from evaluation.
Fine-tuned Encoder-Only Models:
We also evaluate a range of encoder-based models trained end-to-end on each task. Architectures include BERT-base-uncased, RoBERTa-base, DeBERTa-v3-small, ModernBERT-small, and ModernBERT-large. For each model, we append a task-specific MLP head for event occurrence or ordering prediction. The model input is the same flattened prefix of the event sequence used in other settings, tokenized and formatted according to the respective architecture. These models are trained using standard supervised learning objectives and evaluated on the same metrics as other methods: F1 score for event occurrence, and pairwise concordance for temporal ordering. For survival analysis, the final hidden states are passed into a time-to-event prediction head and evaluated using the time-dependent concordance index.
4 Results
Table 1: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from DeepSeek-R1. Bold values indicate best in category amongst all models.
T2S2 (DeepSeek R1) sepsis-10 sepsis-100
F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
LLM-MLP head
DS-L3.3 70B 0.075 0.482 0.796 0.632 0.397 0.578 0.613 0.595
Prompting
L3.3 70B few-shot 0.095 0.380 0.652 0.360 0.452
Encoder-MLP head
ModernBERT-base 0.306 0.576 0.877 0.646 0.428 0.562 0.607 0.558
Encoder-masking-fine-tuned
ModernBERT-base 0.186 0.449 0.697 0.672 0.325 0.576 0.503 0.613
Encoder-masking-zeroshot
ModernBERT-base 0.169 0.468 0.804 0.496 0.358 0.501 0.511 0.499
ModernBERT-large 0.133 0.349 0.534 0.494 0.271 0.509 0.277 0.502
Table 2: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from Llama 3.3-70B. Bold values indicate best in category amongst all models.
T2S2 (Llama 3.3) sepsis-10 sepsis-100
F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
LLM-MLP head
DS-L3.3 70B 0.140 0.563 0.811 0.624 0.397 0.585 0.629 0.598
Prompting
L3.3 70B few-shot ordinal 0.064 0.460 0.729 0.313 0.480
Encoder-MLP head
ModernBERT-base 0.257 0.645 0.903 0.594 0.431 0.590 0.626 0.565
Encoder-masking-fine-tuned
ModernBERT-base 0.139 0.489 0.700 0.632 0.334 0.595 0.478 0.612
Encoder-masking-zeroshot
ModernBERT-base 0.132 0.525 0.840 0.556 0.352 0.503 0.511 0.501
4.1 Event forecast within next 24 hours: F1 performance
Our results demonstrate a clear advantage of encoder-based models over LLMs in event forecasting, for both textual time-series annotations generated from DeepSeek-R1 (Table 1) and Llama-3.3-70B (Table 2) and particularly in short-term and medium-term prediction windows. The complete forecasting performance for all models have been shown in Table 6 and Table 7 in Appendix A. Encoder models, particularly those with a fine-tuned MLP head, achieve substantially higher F1 scores than the LLMs, reinforcing the effectiveness of encoder-based representations for forecasting.
Among encoder models, different training strategies yield distinct performance trends. The MLP head fine-tuning approach consistently outperforms both fine-tuned masking and zero-shot masking models, particularly in long-horizonforecasting. ModernBERT-base achieves the highest overall F1 scores across all forecasting horizons, particularly at 168 hours, where it outperforms all other models. ModernBERT-large follows closely but does not consistently surpass its base variant, suggesting diminishing returns with increased model size. Meanwhile, zero-shot masking models, particularly standard BERT and RoBERTa, fail to make meaningful predictions, often yielding near-zero F1 scores at the 1-hour mark. However, DeBERTa-small and ModernBERT-base exhibit relatively better zero-shot performance, indicating that certain architectures generalize better without task-specific training.
Performance trends also vary across different datasets. Across all models, F1 scores improve as the forecasting window increases, with 1-hour predictions being the most challenging and 168-hour predictions yielding the highest scores. This suggests that longer-term trends are more predictable, whereas immediate event forecasting remains difficult, possibly due to the sparsity and volatility of short-term event occurrences. Comparing internal and external datasets, models perform worse on external validation sets (sepsis-10 and sepsis-100) than on the internal T2S2 test set, with a particularly noticeable drop in F1 at the 1-day horizon. This indicates potential generalization challenges, as models trained on internal data struggle to adapt to distributional shifts in clinical settings. Notably, ModernBERT models maintain a relative performance edge even in external datasets, reinforcing their robustness across different domains.
4.2 Temporal ordering of forecasted events: Concordance
In addition to event occurrence prediction, models are evaluated on their ability to correctly rank the order of upcoming events, measured through concordance (c-index) for both textual time-series annotations generated from DeepSeek-R1 (Table 1) and Llama-3.3-70B (Table 2). The complete forecasting performance for all models have been shown in Table 6 and Table 7 in Appendix A.
Similar to F1 trends, encoder-based models significantly outperform LLMs, with LLMs failing to achieve reliable ordinal rankings even when fine-tuned or prompted. Within encoder models, different training approaches show distinct performance characteristics: MLP-head fine-tuned models dominate in event occurrence classification (F1) but do not always yield the best concordance values. Instead, fine-tuned masking models, particularly ModernBERT-large, achieve the highest concordance scores, suggesting that this approach is better suited for ordering rather than binary classification.
Performance differences across datasets also highlight the challenges of generalization. While internal (T2S2) test results show relatively strong concordance for fine-tuned models, c-index values drop considerably on external datasets, particularly for models that are not ModernBERT-based. Among the best-performing models, ModernBERT-large (fine-tuned masking) achieves the highest c-index across both internal and external datasets, reinforcing its strength in capturing event order relationships. However, zero-shot masking models continue to struggle, mirroring their poor F1 performance. Additionally, while MLP-head fine-tuned models maintain high F1 scores, they exhibit slightly lower concordance than fine-tuned masking models, particularly at the 168-hour horizon. This suggests that different optimization strategies may be required depending on whether the task prioritizes strict classification accuracy or relative event ordering.
Overall, the results confirm that ModernBERT architectures are the strongest performers across both F1 and concordance evaluations, with fine-tuned MLP-head models excelling in binary prediction tasks and fine-tuned masking models achieving superior ranking performance. However, all models face challenges in generalizing to external datasets, emphasizing the need for domain adaptation techniques to improve real-world forecasting robustness.
4.3 Survival analysis
Model Performance across Timepoints and Cohorts.
Table 3 presents the time-dependent concordance index (concordance_td) for all models across three evaluation cohorts—t2s2, Sepsis10, and Sepsis100—at three observation windows (0h, 24h, and 168h). For each cohort and timepoint, we report performance from the best-fitting survival model (selected based on validation concordance).
Across cohorts, we observe that providing additional observation time (e.g., from 0h to 168h) does not consistently improve predictive performance. In the t2s2 cohort, several models achieve their peak concordance at the earliest or intermediate timepoints (e.g., DeepSeek-R1-Llama-70B peaks at 0h, while Llama-3.3-70B-Instruct peaks at 24h), with some degradation at 168h. In Sepsis100, certain models benefit from longer observation (e.g., DeepSeek-R1-Distill-Llama-8B improves steadily across timepoints), but others decline or plateau. The Sepsis10 cohort exhibits high and uniform performance across models, particularly after 24h, suggesting potential ceiling effects or limited evaluation variability in that subset.
Instruction-tuned LLMs such as Llama-3.3-70B-Instruct, DeepSeek-R1-Distill- Llama-70B, and Llama-3.1-8B-Instruct consistently outperform conventional transformer baselines like bert-base-uncased or roberta-base. This trend is particularly notable in the t2s2 and Sepsis100 cohorts, where LLMs demonstrate stronger early signal detection at 0h and 24h. These results suggest that large-scale pretraining and instruction tuning confer advantages in extracting actionable clinical representations, even with limited temporal context.
Taken together, these findings highlight that strong early performance does not require extended observation windows and that pretrained instruction-following models can offer substantial benefits for survival prediction in real-world, noisy clinical text scenarios.
Model t2s2 Sepsis-10 Sepsis-100
0h 24h 168h 0h 24h 168h 0h 24h 168h
bert-base-uncased 0.60 0.61 0.53 1.00 1.00 1.00 0.58 0.62 0.58
roberta-base 0.55 0.60 0.60 0.75 1.00 1.00 0.55 0.58 0.57
deberta-v3-small 0.56 0.57 0.57 0.75 0.00 0.75 0.52 0.49 0.51
ModernBERT-base 0.52 0.58 0.58 1.00 1.00 1.00 0.54 0.51 0.51
ModernBERT-large 0.53 0.59 0.58 0.50 1.00 1.00 0.59 0.55 0.49
DeepSeek-R1-Llama-70B 0.64 0.63 0.59 0.50 1.00 1.00 0.61 0.53 0.55
Llama-3.3-70B-Instruct 0.62 0.63 0.58 0.50 1.00 0.75 0.64 0.57 0.53
DeepSeek-R1-Llama-8B 0.60 0.58 0.58 1.00 1.00 1.00 0.53 0.66 0.65
Llama-3.1-8B-Instruct 0.62 0.61 0.61 1.00 1.00 1.00 0.60 0.59 0.61
Table 3: Concordance index (concordance_td) for each model across three test sets (t2s2, Sepsis10, Sepsis100), evaluated at three observation time thresholds (0h, 24h, 168h).
4.4 Sensitivity analysis
To better understand the robustness and design implications of our forecasting models, we conduct two complementary sensitivity analyses focused on the temporal structure and completeness of input sequences. First, we compare two strategies for ordering clinical events within textual time-series (TTS) inputs: text-ordered, where events follow their extraction sequence from the original case report narratives, and time-ordered, where events are chronologically sorted. This comparison assesses the relative benefits of preserving narrative flow versus enforcing temporal alignment. Second, we evaluate the impact of input sparsity by randomly masking portions of the clinical history, simulating scenarios with incomplete documentation or missing data. Together, these analyses shed light on the importance of chronological ordering and data completeness for both classification accuracy and temporal event ranking.
4.4.1 Time-ordered vs text-ordered training
We evaluated the impact of training on text-ordered versus time-ordered textual time-series (TTS) annotations. In the text-ordered condition, events in the textual time-series order follow the order in which they were extracted from the case reports using LLMs (DeepSeek-R1 and Llama-3.3-70B), while in the time-ordered condition, they are sorted chronologically in time. This comparison aimed to assess the impact of chronological versus narrative ordering of events on forecasting performance
As shown in Table 4, time-ordered training generally leads to higher concordance (C-index), particularly for DeepSeek-R1 annotatiobs across both internal (T2S2) and external test datasets (sepsis-10 and sepsis-100). However, F1 scores are less consistently affected by the ordering strategy. In several cases—especially with Llama-3.3-70B annotations on external datasets, text-ordered training results in equal or better F1 scores, suggesting that preserving the narrative sequence can support accurate classification even when it does not follow temporal order. These results highlight a trade-off between chronological alignment and narrative structure in optimizing forecasting performance.
Table 4: Comparison of text-ordered and time-ordered training for Deep Seek R1 and Llama 3.3 annotations on F1 scores and C-index values.
Dataset Metric DeepSeek-R1 Llama 3.3-70B
Text-ordered Time-ordered Text-ordered Time-ordered
T2S2 Test F1 (1hr) 0.290 0.306 0.311 0.257
F1 (1day) 0.569 0.576 0.648 0.645
F1 (7days) 0.876 0.877 0.898 0.903
C-index 0.659 0.672 0.586 0.632
Sepsis10 F1 (1day) 0.361 0.428 0.423 0.382
C-index 0.598 0.576 0.526 0.595
Sepsis100 F1 (1day) 0.615 0.607 0.662 0.616
C-index 0.609 0.613 0.541 0.612
4.4.2 Impact of Randomly Masking History on F1 and Concordance
Further, we systematically evaluate how randomly omitting portions of clinical history impacts both forecasting accuracy and the model’s ability to correctly order clinical events. Towards this end, we simulate incomplete patient histories by randomly masking input time steps at varying rates. We refer to this as the timestep drop rate (TDR), ranging from 0% (no masking) to 90% (severely truncated history). We assess the effect of this masking on both classification performance (F1 score at 1 day) and ranking performance (concordance index) across three datasets—T2S2-DS, Sepsis-10, and Sepsis-100—using both DeepSeek-R1 (Table 5) and Llama-3.3-70B annotations (Table 8 in Appendix B).
F1 scores decline progressively with increasing timestep dropout rate (TDR), with sharp reductions beyond 60% (Table 5). At TDR = 0%, the F1 scores for T2S2-DS, Sepsis-10, and Sepsis-100 are 0.576, 0.428, and 0.607, respectively. By TDR = 90%, these values fall to 0.423, 0.200, and 0.510, with Sepsis10 showing the steepest decline. Between 30% and 60% dropout, performance remains relatively stable for T2S2-DS and Sepsis100, while Sepsis10 exhibits a more pronounced drop at TDR = 60%. Similar trends are observed with Llama-3.3 annotations (Table 8 in Appendix B).
In contrast, concordance (c-index) remains largely consistent across dropout levels. T2S2-DS (Table 5) reaches its highest concordance (0.678) at TDR = 60%, while Sepsis-10 and Sepsis-100 vary only marginally. This suggests that while masking temporal history substantially reduces classification accuracy, it has limited impact on event ranking performance. Comparable patterns are observed for the Llama-3.3 models (Table 8).
Overall, DeepSeek models show notable degradation in F1 scores when more than 30% of history is removed—particularly for datasets sensitive to longer-term dependencies—whereas concordance scores remain relatively stable (Table 5 and Table 8).
Table 5: Effect of Randomly Masking History (DeepSeek R1 annotations) on F1 and Concordance. TDR : Timestep drop rate
TDR F1 (1 day) Concordance (c-index)
T2S2-DS Sepsis-10 Sepsis-100 T2S2-DS Sepsis-10 Sepsis-100
0% 0.576 0.428 0.607 0.672 0.576 0.613
15% 0.572 0.415 0.604 0.667 0.564 0.616
30% 0.548 0.357 0.591 0.672 0.542 0.619
60% 0.542 0.310 0.588 0.678 0.578 0.616
90% 0.423 0.200 0.510 0.661 0.561 0.614
5 Discussion and Conclusion
Our findings highlight the superiority of encoder-based models over decoder-based LLMs for event forecasting, emphasizing the limitations of autoregressive models in structured prediction tasks. Despite prompting and fine-tuning, LLMs struggle with short- and medium-term forecasting, reinforcing the need for models optimized for sequence encoding rather than generative token prediction.
Among encoder models, ModernBERT-base and ModernBERT-large consistently perform best, with MLP-head fine-tuning excelling in F1 scores and fine-tuned masking models achieving higher concordance (c-index), particularly in long-term forecasting. This suggests that classification and ranking tasks benefit from distinct optimization strategies, and a hybrid training approach may further enhance performance.
The gradual improvement of F1 scores over longer time horizons indicates that event patterns become more predictable over time, whereas short-term forecasting remains challenging due to higher variability. This limitation suggests that incorporating temporal modeling techniques, such as recurrent or convolutional structures, could refine short-term predictions while maintaining strong long-range forecasting.
Performance drops on external validation datasets (sepsis-10 and sepsis-100) highlight generalization challenges, with models struggling to adapt to shifts in event distributions. While ModernBERT models exhibit relative robustness, domain-specific fine-tuning or adaptive learning strategies are needed to bridge this gap.
Our survival analysis results complement the forecasting findings by revealing that strong predictive performance is often achievable with limited observation windows. Several models reach peak concordance at early (0h) or intermediate (24h) timepoints rather than at the longest window (168h), particularly in the T2S2 and Sepsis100 cohorts. Instruction-tuned LLMs such as Llama-3.3-70B-Instruct and DeepSeek-R1-Distill-Llama-70B outperform traditional transformer baselines in these survival settings, suggesting that instruction tuning and large-scale pretraining confer meaningful advantages for modeling time-to-event outcomes from clinical narratives in the textual time series setup.
Sensitivity analyses further highlight the importance of temporal structure and data completeness. When comparing time-ordered and text-ordered training regimes, we find that time-ordered training generally improves concordance, particularly for DeepSeek-R1 annotations, while text-ordering can yield competitive or better F1 scores on external datasets. This underscores a trade-off between temporal alignment and narrative coherence, with implications for designing learning objectives based on downstream task priorities. Moreover, robustness experiments simulating documentation sparsity reveal that F1 scores degrade significantly as more clinical history is masked, particularly beyond a timestep dropout rate of 60%. In contrast, concordance remains relatively stable under these conditions, indicating that event ranking is less sensitive to partial history than event classification.
Overall, our results demonstrate that encoder-based models, particularly ModernBERT, are well-suited for event forecasting but require enhancements in short-term prediction and cross-domain generalization. In contrast, for survival analysis tasks, we find that instruction-tuned LLMs have more predictive power. Future work should explore hybrid training objectives that balance classification and ranking performance, incorporate temporal modeling architectures, and investigate domain adaptation and pretraining strategies that explicitly account for data sparsity and real-world variability that we observe in our clinical textual time series.
Acknowledgments
This research was supported in part by the Division of Intramural Research (DIR) of the National Library of Medicine (NLM), National Institutes of Health. This work utilized the computational resources of the NIH HPC Biowulf cluster.
Appendix A Forecasting performance on DeepSeek-R1 and Llama-3.3-70B textual time-series annotations
Tables 6 and 7 present the forecasting performance of our models on textual time-series annotations generated by DeepSeek-R1 and Llama-3.3-70B, respectively. We report standard forecasting metrics across multiple evaluation settings to provide a comprehensive view of model performance.
Table 6: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from DeepSeek-R1. Bold indicates best in that particular category amongst all models.
T2S2 (DeepSeek R1) sepsis-10 sepsis-100
F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
LLM-MLP head
DS-L3.3 70B 0.075 0.482 0.796 0.632 0.397 0.578 0.613 0.595
L3.3 70B 0.000 0.295 0.764 0.631 0.312 0.569 0.404 0.597
DS-L3.1 8B 0.000 0.492 0.650 0.630 0.381 0.566 0.601 0.600
L3.1 8B 0.000 0.355 0.650 0.630 0.114 0.575 0.442 0.604
Prompting
L3.3 70B 0-shot 0.085 0.368 0.560 0.236 0.452
L3.3 70B 0-shot ordinal 0.082 0.352 0.646 0.289 0.483
L3.3 70B few-shot 0.095 0.380 0.652 0.360 0.452
L3.3 70B few-shot ordinal 0.063 0.356 0.682 0.313 0.480
Encoder-MLP head
BERT 0.264 0.550 0.871 0.501 0.441 0.498 0.614 0.501
RoBERTa 0.250 0.562 0.866 0.500 0.405 0.500 0.614 0.500
DeBERTa-small 0.262 0.570 0.871 0.601 0.382 0.498 0.617 0.556
ModernBERT-base 0.306 0.576 0.877 0.646 0.428 0.562 0.607 0.558
ModernBERT-large 0.277 0.568 0.874 0.598 0.391 0.539 0.602 0.553
Encoder-masking-fine-tuned
BERT 0.166 0.432 0.690 0.648 0.323 0.554 0.482 0.600
RoBERTa 0.169 0.416 0.688 0.646 0.313 0.539 0.457 0.609
DeBERTa-small 0.175 0.435 0.695 0.671 0.325 0.579 0.492 0.617
ModernBERT-base 0.186 0.449 0.697 0.672 0.325 0.576 0.503 0.613
ModernBERT-large 0.177 0.460 0.683 0.663 0.342 0.615 0.496 0.632
Encoder-masking-zeroshot
BERT 0.000 0.000 0.000 0.495 0.000 0.513 0.000 0.494
RoBERTa 0.049 0.207 0.375 0.500 0.137 0.500 0.318 0.500
DeBERTa-small 0.170 0.460 0.323 0.497 0.268 0.512 0.376 0.509
ModernBERT-base 0.169 0.468 0.804 0.496 0.358 0.501 0.511 0.499
ModernBERT-large 0.133 0.349 0.534 0.494 0.271 0.509 0.277 0.502
Table 7: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from Llama-3.3-70B.
T2S2 (Llama 3.3) sepsis-10 sepsis-100
F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
LLM-MLP head
DS-L3.3 70B 0.140 0.563 0.811 0.624 0.397 0.585 0.629 0.598
L3.3 70B 0.265 0.466 0.807 0.619 0.160 0.567 0.466 0.597
DS-L3.1 8B 0.264 0.584 0.837 0.624 0.258 0.576 0.605 0.598
L3.1 8B 0.338 0.565 0.823 0.619 0.368 0.569 0.536 0.596
Prompting
L3.3 70B 0-shot 0.085 0.443 0.616 0.236 0.452
L3.3 70B 0-shot ordinal 0.063 0.454 0.707 0.289 0.483
L3.3 70B few-shot 0.099 0.457 0.703 0.360 0.452
L3.3 70B few-shot ordinal 0.064 0.460 0.729 0.313 0.480
Encoder-MLP head
BERT 0.224 0.622 0.899 0.500 0.400 0.500 0.612 0.500
RoBERTa 0.180 0.608 0.900 0.500 0.364 0.500 0.597 0.500
DeBERTa-small 0.258 0.641 0.898 0.603 0.427 0.585 0.616 0.592
ModernBERT-base 0.257 0.645 0.903 0.594 0.431 0.590 0.626 0.565
ModernBERT-large 0.312 0.622 0.897 0.500 0.404 0.500 0.623 0.500
Encoder-masking-fine-tuned
BERT 0.131 0.471 0.741 0.622 0.320 0.580 0.467 0.622
RoBERTa 0.129 0.484 0.659 0.625 0.329 0.587 0.472 0.614
DeBERTa-small 0.135 0.468 0.693 0.648 0.318 0.602 0.462 0.623
ModernBERT-base 0.139 0.489 0.700 0.632 0.334 0.595 0.478 0.612
ModernBERT-large 0.124 0.470 0.712 0.646 0.318 0.586 0.468 0.631
Encoder-masking-zeroshot
BERT 0.000 0.000 0.000 0.504 0.000 0.515 0.000 0.506
RoBERTa 0.074 0.279 0.445 0.500 0.083 0.500 0.319 0.500
DeBERTa-small 0.008 0.439 0.839 0.562 0.296 0.518 0.518 0.501
ModernBERT-base 0.132 0.525 0.840 0.556 0.352 0.503 0.511 0.501
ModernBERT-large 0.109 0.413 0.622 0.499 0.326 0.509 0.339 0.518
Appendix B Masking history: Llama-3.3-70B textual time-series annotations
Table 8 examines the impact of randomly masking portions of the input history on model performance when using Llama-3.3-70B-generated annotations. We evaluate performance across three datasets (T2S2-L33, Sepsis-10, and Sepsis-100) using F1 scores for prediction of next $k$ events happening with the next day and concordance index (c-index) correctly ordering the next $k$ events. As the proportion of masked history increases, we observe a general decline in F1 performance, particularly at higher masking rates (e.g., 90%), suggesting that access to historical context is important for accurate forecasting. Interestingly, the concordance index remains relatively stable, indicating the model’s robustness in ranking the next $k$ events with partial input histories.
Table 8: Effect of Randomly Masking History on F1 and Concordance (Llama-3.3 Annotations)
TDR F1 (1 day) Concordance (c-index)
T2S2-L33 Sepsis-10 Sepsis-100 T2S2-L33 Sepsis-10 Sepsis-100
0% 0.645 0.431 0.626 0.632 0.595 0.612
15% 0.628 0.422 0.624 0.635 0.592 0.610
30% 0.634 0.423 0.621 0.634 0.586 0.598
60% 0.604 0.391 0.605 0.634 0.577 0.613
90% 0.531 0.328 0.534 0.622 0.546 0.607
Appendix C Textual Time Series Forecasting Framework of Encoder Models
We develop a general framework for predictive modeling over textual time series of the encoder-only models, designed to accommodate the irregular, event-based structure common to clinical case reports. As highlighted Section 3.2, each data instance consists of a sequence of clinical events, where each event is a free-text string describing a clinical finding, intervention, or diagnosis, paired with a timestamp in hours relative to the time of admission. The format resembles a two-column table, where the first column contains textual clinical events and the second column contains integer-valued timestamps.
Input Preprocessing and Timeseries Construction.
From each case report, we construct multiple training examples for sequence forecasting. For a fixed forecast horizon parameter $K$, we generate overlapping windows of historical events and forecast targets. Specifically, for each case report with $L$ rows, we construct a sequence of up to $L-K$ forecasting examples. Each example is structured as follows:
  • A history segment $\mathcal{H}_{t}=\{(e_{i},t_{i})\}_{i=1}^{n}$, consisting of clinical events $e_{i}$ occurring at timestamps $t_{i}\leq t$.
  • A forecast target $\mathcal{F}_{t}=\{(f_{j},\tau_{j})\}_{j=1}^{K}$, consisting of the next $K$ events after $t$.
To preserve the temporal coherence of events, we treat all events with the same timestamp as an atomic unit. That is, if multiple events share the same timestamp, they are included together in the history or forecast block, and the next forecasting window is only advanced after the full timestamp block has been consumed.
Model Input Representation.
Each training example is serialized into a transformer-compatible input by linearizing the historical events and forecast events. Importantly, we preserve both temporal and textual content by encoding each event in the form “t:e” where t is the timestamp and e is the event string. The input to the language model takes the form:
[…]
This is followed by additional target-specific components, as described below.
Language Models.
To test the generality of our framework, we experiment with a variety of pre-trained encoder-only language models, including:
  • bert-base-uncased (BERT; Devlin et al., 2018)
  • roberta-base (Liu et al., 2019)
  • microsoft/deberta-v3-small (He et al., 2021)
  • answerdotai/ModernBERT-base and answerdotai/ModernBERT-large (pretrained on modern clinical narratives)
Each model is used with its corresponding tokenizer. For models with SentencePiece tokenization (e.g., DeBERTa), we use the appropriate fast tokenizer such as DebertaV2Tokenizer to ensure compatibility. All inputs are truncated from the left (i.e., keeping the most recent tokens) if they exceed the model’s context window (typically 512 tokens).
Prediction Tasks.
We define two primary prediction tasks over $\mathcal{F}_{t}$, each formulated as a binary classification problem.
1. Event Ordering (Concordance) Task.
The goal in this task is to predict the correct temporal ordering between future events. For each example, we generate all comparable pairs of events in $\mathcal{F}_{t}$, defined as pairs $(f_{i},f_{j})$ where $\tau_{i}\neq\tau_{j}$. For each pair, we form a binary classification example as follows:
  • Input: The serialized history segment followed by:
[…]
  • Label: $y=1$ if $\tau_{i}<\tau_{j}$, else $y=0$
The model output is taken from the [CLS] token, passed through a linear classification head, and softmaxed to produce a binary probability. We optimize cross-entropy loss across all pairwise comparisons. At evaluation time, we report the concordance index (c-index), which measures the proportion of correctly predicted pairs among all comparable pairs.
2. Time-Window Classification Task.
This task evaluates whether each forecast event occurs within a fixed prediction window $H$ hours of the last historical event. For each $f_{j}\in\mathcal{F}_{t}$, we generate an input:
  • Input: The serialized history segment followed by:
[…]
  • Label: $y=1$ if $\tau_{j}-t_{n}\leq H$, where $t_{n}$ is the timestamp of the last historical event.
This is again treated as a binary classification task over the [CLS] embedding, using a linear layer and softmax. We report macro-averaged F1 scores over all prediction examples.
Training and Evaluation.
Each task is trained separately using AdamW with linear learning rate decay and early stopping based on validation loss. We save the model checkpoint with the best validation performance. During inference, we reload this checkpoint and evaluate on a held-out test set (which is obtained from textual time series of unseen case reports).
For both tasks, we extract softmax probabilities rather than hard labels to allow for threshold calibration and ROC-based analyses. All training is conducted using PyTorch and HuggingFace Transformers.
Appendix D Survival Modeling with Language Model Embeddings
To evaluate the prognostic value of textual information encoded in large language models (LLMs), we adopt a two-stage framework: (1) extraction of fixed-dimensional sequence embeddings from various pre-trained LLMs, and (2) downstream survival modeling using these embeddings as covariates.
In the first stage, we process each textual time series (which is converted to a textual context of ”(time) clinical event [SEP] … [SEP] (time) clinical event [SEP]”) using a suite of LLMs to obtain dense vector representations that summarize the content of the input sequence. For models belonging to the encoder family, including bert-base-uncased, roberta-base, deberta-v3-small, ModernBERT-base, and ModernBERT-large, we extract the final hidden state corresponding to the [CLS] token, which is conventionally used to represent the entire sequence in classification tasks. This token-specific embedding serves as a compact, sequence-level representation.
For decoder-based models that do not utilize a dedicated [CLS] token, including DeepSeek-R1-Distill-Llama-70B, Llama-3.3-70B-Instruct, DeepSeek-R1-Distill-Llama-8B, and Llama-3.1-8B-Instruct, we compute the mean-pooled embedding over the last hidden states of all non-padding tokens. This pooling strategy yields a fixed-length vector that captures the overall semantic content of the input while mitigating the impact of padding artifacts.
In the second stage, these embedding vectors are used as input covariates to three survival models, each designed to capture time-to-event dynamics in different ways:
  • Random Survival Forest (RSF): A nonparametric ensemble method based on decision trees, capable of handling complex interactions and high-dimensional inputs. The model is tuned over the number of trees (n_estimators), the minimum number of samples required to split an internal node (min_samples_split), and the minimum number of samples at a leaf node (min_samples_leaf).
  • DeepSurv: A neural-network generalization of the Cox proportional hazards model. The architecture comprises fully connected layers with varying widths (num_nodes) and dropout rates, with a fixed number of training epochs.
  • DeepHit: A multi-task neural network that jointly models the discrete hazard and survival functions. As with DeepSurv, we perform a grid search over hidden layer sizes and dropout rates while keeping the number of epochs fixed.
Hyperparameter tuning for each survival model is performed using a validation split from the training set. For each combination of language model and survival model, we conduct grid search over the respective hyperparameter space, selecting the configuration that maximizes the time-dependent concordance index on the validation set. Final evaluation is conducted on held-out test sets (t2s2_test, sepsis10, sepsis100), and results are aggregated across survival metrics to assess model performance.
This framework allows for a systematic evaluation of how different pre-trained LLMs, and their associated embedding strategies, contribute to downstream survival prediction tasks.
Appendix E LLM prompt to generate clinical textual time-series from PMOA case reports
You are a physician. Extract the clinical events and the related time stamp from the case report. The admission event has timestamp 0. If the event is not available, we treat the event, e.g. current main clinical diagnosis or treatment with timestamp 0. The events that happened before event with 0 timestamp have negative time, the ones after the event with 0 timestamp have positive time. The timestamp are in hours. The unit will be omitted when output the result. If there is no temporal information of the event, please use your knowledge and events with temporal expression before and after the events to provide an approximation. We want to predict the future events given the events happened in history. For example, here is the case report.
An 18-year-old male was admitted to the hospital with a 3-day history of fever and rash. Four weeks ago, he was diagnosed with acne and received the treatment with minocycline, 100 mg daily, for 3 weeks. With increased WBC count, eosinophilia, and systemic involvement, this patient was diagnosed with DRESS syndrome. The fever and rash persisted through admission, and diffuse erythematous or maculopapular eruption with pruritus was present. One day later the patient was discharged.
Let’s find the locations of event in the case report, it shows that four weeks ago of fever and rash, four weeks ago, he was diagnosed with acne and receive treatment. So the event of fever and rash happened four weeks ago, 672 hours, it is before admitted to the hospital, so the time stamp is -672. diffuse erythematous or maculopapular eruption with pruritus was documented on the admission exam, so the timestamp is 0 hours, since it happens right at admission. DRESS syndrome has no specific time, but it should happen soon after admission to the hospital, so we use our clinical judgment to give the diagnosis of DRESS syndrome the timestamp 0. then the output should look like: 18 years old | 0
male | 0
admitted to the hospital | 0 fever | -72
rash | -72
acne | -672
minocycline | -672
increased WBC count | 0
eosinophilia| 0
systemic involvement| 0
diffuse erythematous or maculopapular eruption| 0
pruritis | 0
DRESS syndrome | 0
fever persisted | 0
rash persisted | 0
discharged | 24
Separate conjunctive phrases into its component events and assign them the same timestamp (for example, separation of ‘fever and rash’ into 2 events: ‘fever’ and ‘rash’). If the event has duration, assign the event time as the start of the time interval. Attempt to use the text span without modifications except ‘history of’ where applicable. Include all patient events, even if they appear in the discussion; do not omit any events; include termination/discontinuation events; include the pertinent negative findings, like ‘no shortness of breath’ and ‘denies chest pain’. Show the events and timestamps in rows, each row has two columns: one column for the event, the other column for the timestamp. The time is a numeric value in hour unit. The two columns are separated by a pipe ‘|’ as a bar-separated file. Skip the title of the table. Reply with the table only.
Appendix F Prompting strategies for forecasting tasks
Zero-shot:
You are an expert physician.
Reply to the prompt with structured predictions in a k-item, bar-separated row. For example, if there k=3 events (A, B, C) and only B occurs in the time window, then: 0 | 1 | 0 would be the correct response. Reply with the answer only. Do NOT provide any other text. Do NOT explain.
Zero-shot ordinal:
You are an expert physician.
Reply to the prompt with structured predictions in a k-item, bar-separated row. Assess whether the event happens in the time window using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen. For example, if there k=3 events (A, B, C) and only B occurs in the time window, then: $1|5|1$ would be the correct response. Reply with the answer only. Do NOT provide any other text. Do NOT explain.
Few-shot:
You are an expert physician.
Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window. Example 1. If there k=3 events (A, B, C) and only B occurs in the requested time window, then: 0 | 1 | 0 would be the correct response.
Example 2. Suppose the time series is:
admitted | 0
history of diabetes | 0
elevated troponin | 5
--- PREDICTION WINDOW ---
cardiac catheterization | 6
admission to icu | 24
transfer to floor | 72
discharge | 144
office visit | 1440
Then, if the prediction is at 5 hours, the time preceding ‘‘--- PREDICTION WINDOW ---’’, the events admitted, history of diabetes, and elevated troponin have occurred. Supposing k=4 and a forecast window of 24 hours, then the task it to determine which of cardiac catheterization, admission to icu, transfer to floor, and discharge happen. So the correct response for the example is: 1 | 1 | 0 | 0 because cardiac catherization (6) and admission to icu (24) happen before 5 + 24 = 29.
Reply with the answer only. Do NOT provide any other text. Do NOT explain.
Few-shot ordinal:
You are an expert physician.
Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window, using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen.
Example 1. If there k=3 events (A, B, C) and only B occurs in the time window, then: 1 | 5 | 1 would be the correct response.
Example 2. Suppose the time series is:
admitted | 0
history of diabetes | 0
elevated troponin | 5
--- PREDICTION WINDOW ---
cardiac catheterization | 6
admission to icu | 24
transfer to floor | 72
discharge | 144
office visit | 1440
Then, if the prediction is at 5 hours, the time preceding ‘‘--- PREDICTION WINDOW ---’’, the events admitted, history of diabetes, and elevated troponin have occurred. Supposing k=4 and a forecast window of 24 hours, then the task it to determine which of cardiac catheterization, admission to icu, transfer to floor and discharge happen. So the an excellent response for the example is: 5 | 4 | 1 | 1 because cardiac catherization (6) and admission to icu (24) happen before 5 + 24 = 29.
Reply with the answer only. Do NOT provide any other text. Do NOT explain.